OpenStudy (anonymous):

“The length of the second side of a triangle is one foot less than three times the length of the first side. The length of the third side is two feet more than the length of the first side. If the perimeter of the triangle must be larger than 108 feet, what are the dimensions of the triangle

OpenStudy (anonymous):

if x feet is the length of the first side, what can you tell me about the lengths of the second and third side?

OpenStudy (anonymous):

i need your feedback to help you, bella

OpenStudy (anonymous):

sorry i was working on another problem

OpenStudy (anonymous):

Let x be the length of the first side. 3x-1 be the length of the second side x+2 be the length of the third side. P=108 P=a+b+c 108=x+3x-1+x+2 108=5x+1 5x=108-1 5x=107 x=21.4 ft(length of first side) 3x-1=3(21.4)-1 =64.2-1 =63.2(length of second side) x+2=21.4+2=23.4(length of the third side)

OpenStudy (anonymous):

omg pat18 tks this is awesome i was having such a hard time figuring this one out

OpenStudy (anonymous):

thanks. i hope this helps.

OpenStudy (anonymous):

i will find out when i actually do the actual math problem on friday

OpenStudy (anonymous):

okay.thank you

OpenStudy (anonymous):

perimeter is GREATER than 108, so x > 21.4 and so on

OpenStudy (anonymous):

yeah. you can just adjust it.

OpenStudy (anonymous):

tks pat

OpenStudy (anonymous):

Say x is the length of first side den 3x-1 and x+2 are the 2nd and 3rd sides respectively ..by triangle inequality x+x+2>3x-1 => x<3 .. perimeter is 5x+1 which <16 for x < 3 .. so such a triangle cant b possible

OpenStudy (anonymous):

you are welcome bellarue. it is my pleasure to help. ^_^

OpenStudy (anonymous):

23.4+21.4=44.8 which s less than ur third side 63.2 ... for a triangle formation the sum o oder two sides should be less than d third side ... Basic definition

OpenStudy (anonymous):

that is why the problem states it is greater than 21.4 and so on.

OpenStudy (anonymous):

it is not actually 23.4 + 21.4

OpenStudy (anonymous):

Dude all the three sides are dependent on x .. See x+x+2> 3x-1 ...(applying the triangle side rule ..) => x< 3 for this inequality .... which s contradictin to ur x > 21.4

OpenStudy (anonymous):

yup, dk is right